Sol Summary – April 17th

Crew 297 Sol Summary Report 17-04-2024

Sol: 03

Summary Title: Red Planet Rocks!

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary: The crew started the day at 0800, welcomed by the smell of the delectable crepes prepared by Sean and Matt S. With the mood already in good spirits after breakfast, the EVA team could not flounder, as they hastily were able to locate the NPS and spent the rest of the time analyzing the rocks of the Red Planet. After a fortunate EVA, the crew was treated to a mouthwatering take on fromage et legumes soup prepared by Matt L. While the soup was prepared, the crew learned how to play the cooperative card game Hanabi. In the afternoon, Sean and Pawel troubleshooted thermal problems within the GreenHab along with harvesting a ripe cherry tomato. Dinner concluded a day of scrumptious meals, with a soy peanut couscous containing vegetables and chicken, prepared by chefs Matt L. and Sean.

Look Ahead Plan:

Tomorrow we will perform a single EVA in the morning. This EVA will be the first for Case, our robotic crewmate, who will traverse and record imagery as we monitor closely. The EVA will also test out digging possibilities for the [inert] NPS.

Anomalies in work:

Suit 5 (ventilation electrical connection): Investigated and likely fixed. To be further examined/tested in the coming days.
GreenHab automated fan: The GreenHab fan powered on today when the GH temp reached 94F, but then turned off. It then powered cycled continuously but never reached its expected rpm.
Suit 4 (collar cable ring missing): In work
Morning (data from 0900): calm winds (2 mph), chilly (53F), and clear skies (0% cloud coverage)
Noon: moderate winds (15 mph), warm (70F), and clear skies (0% cloud coverage)
Afternoon (1800): moderate winds (17 mph), warm (76F), and clear skies (0% cloud coverage)
Crew Physical Status: Crew health is nominal.

EVA: EVA was nominal today. The crew found the hidden NPS fairly quickly (~15 minutes) and spent the rest of the time performing gamma-ray spectrometer readings of regions along Cow Dung Road south of MDRS including Robert’s Rock Garden.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (1 EVA)
EVA request (1 EVA)
Support Requested: We request that the [inert] NPS be placed by Mission Support at a location of their choosing (though reachable by rover and within a maximum of 15 minutes of walking) in either Candor Chasma, Eos Chasma, Valles Marineris, or Aurorae Chaos for a search and recovery EVA on 19-04-2024. The NPS will be placed in the rear airlock tomorrow night, 18-04-2024 by 1600. Once the inert NPS is hidden, precise GPS coordinates and a photo are requested to be taken to not lose track of the location. These GPS coordinates are then asked to be sent to Pawel.

Copyright © The Mars Society. All rights reserved. | Main Site